OpenStudy (anonymous):

X^2+2X-3=0 im suppose to use a graph so any help will work

OpenStudy (anonymous):

you might want to find the zeros and the vertex. zeros will show you where it crosses the x - axis, and vertex is where this parabola sits.

OpenStudy (anonymous):

i have the graph i just dont know how to draw it here, i tried once and it looked horrible lol

OpenStudy (anonymous):

to find the zeros you can factor as \[x^2+2x-3=(x-1)(x+3)=0\] and so \[x-1=0 \text { or } x+3=0\] \[x=1\text { or } x=-3\] and vertex is at (-1,-4)

OpenStudy (anonymous):

im suppose to use the graph to solve

OpenStudy (anonymous):